### v1

@@ 2.1 Data-Driven Motion Modelling Methods
Contemporary research in motion data modelling predominantly explore learned data-driven methods, using neural networks to capture, model, and establish correlations between the motion features of various skeletal configurations [49, 53, 39]. Motion features alone have demonstrated considerable efficacy in cross-skeleton imitation using a variety of strategies, including task-driven objectives [43, 19], diffusion-based generation [47, 25, 50], and latent feature consistency [2, 1, 44, 3, 7]. Latent consistency in particular aims to learn inter-skeleton correlations without the need for paired datasets, which reduces the barrier of entry for practical applications.
@@ 3.1 Quaternion-based Motion Representations
where $QR$ is the quaternion rotation function, and $\mathbf{v}_{m}$ is the parent joint. In summary, a motion sequence $M$ of $|M|$ frames is defined by:
@@ 3.3 First-Frame Offset Quaternions
Crucially, FOQ is a roll-invariant rotation representation, that is, FOQ remains identical regardless of a bone’s absolute roll position in the original data. Trivially, to adjust the rotational roll of a bone (e.g., $\mathbf{v}_{n}$), we perform the quaternion multiplication for each frame (including the initial frame): $\mathbf{q}_{n,t}\times\mathbf{r}$, where $\mathbf{r}$ is a roll quaternion. As a quaternion on the roll axis of $\mathbf{v}_{n}$, $\mathbf{r}$ is constrained by $\mathbf{r}=\alpha+\beta(x\mathbf{i}+y\mathbf{j}+z\mathbf{k})$, where $x$,$y$,$z$ are the 3D XYZ values of $\mathbf{T}_{n}$, and $\alpha,\beta$ are scalars that control the roll magnitude. Note that $\mathbf{r}$ is constant throughout the sequence. We can then deduce that for any constant $\mathbf{r}$,
@@ 3.4 Motion Interpolation Transformer and Rest Pose Augmentation
To leverage our point cloud-based representation learning for cross-skeleton interpolation, a transformer model - CITL [33] is adapted, allowing for an efficient interpolation pipeline with existing datasets to alternative skeletons. We introduce two key modifications to CITL and its training strategy, focusing on enhancing motion quality. First, quaternion predictions are substituted with FOQ predictions. This is inspired by the RNN-based method $\text{TG}_{complete}$ [18], which predicts quaternion offsets from the preceding frame rather than direct quaternion values, thereby achieving greater generalizability and superior quality in interpolation.
@@
4:$\eta\leftarrow\frac{\mathbf{T}}{|\mathbf{T}|}$triangleright 3D unit coordinates   $[\eta_{x},\eta_{y},\eta_{z}]$
5:$\omega\leftarrow\frac{\mathbf{p}_{\text{IK}}}{|\mathbf{p}_{\text{IK}}|}$triangleright 3D unit coordinates $[\omega_{x},\omega_{y},\omega_{z}]$
6:$\mathbf{q}_{\text{min}}\leftarrow\frac{0+(\eta_{x}+\omega_{x})\mathbf{i}+(\eta_{y}+\omega_{y})\mathbf{j}+(\eta_{z}+\omega_{z})\mathbf{k}}{2}$
7:$\mathbf{q}_{\text{max}}\leftarrow\mathbf{q}_{\text{min}}\times(0+\eta_{x}\mathbf{i}+\eta_{y}\mathbf{j}+\eta_{z}\mathbf{k})$
@@ 4.1 Datasets
  • Human3.6M [22, 6] comprises 3.6 million motion frames. The motions captured in Human3.6M tend to be less dynamic, as they were recorded on a relatively compact 4x3m stage. Its native skeleton includes 5 spinal bones, 4 bones for each limb, and an additional finger bone for each hand.
  • CMU Mocap [10] is a diverse motion dataset encompassing 113 overlapping categories. It features a mix of static and dynamic motions, captured on a spacious 3x8m stage. The CMU skeleton is identical to that of Human3.6M, with the exception of the pelvis, which is divided into two symmetrical hip bones and a lower back bone.
For PC-MRL, we designated the CMU MoCap skeleton as $S_{B}$. Its native dataset is exclusively utilized for testing purposes; while LaFAN1 and Human3.6M are used simultaneously for training.
@@ 4.2 Implementation Details
In our experiments, point cloud obfuscation adopted a 256-point setting for the sampling process with $\sigma=0.05$ and each point was characterized by a 64-dimensional vector. To maintain consistency, we ensure that all bone offsets and motion positions are defined in metre units. For skeletal reconstruction, 4 Point Transformer [52] layers was utilized to construct the neural network. Within each layer, the feature size was doubled, and the size of the point set was downsampled by a factor of 4 through a farthest-first traversal approach. This processing across layers resulted in a singular feature vector. In the final step, two linear layers with ReLU activations were utilized to transform the resulted feature vector in a frame-wise manner, producing the output skeletal representation. During training, we set $K=8$ for the KNN loss to optimize the network. Additionally, the use of unit quaternion values in practice relies on modulus division for constraint, a process that can potentially lead to exploding gradients. To address this, we introduced an objective function $\mathcal{L}_{q}$ that measures the $\ell_{2}$ norms’ difference between the raw quaternion output and the expected unit quaternion. All experiments were trained and evaluated on a single NVIDIA GeForce RTX 3090 GPU.
@@ 4.3 Motion Interpolation Comparison against Supervised Methods
To demonstrate the efficacy of our method, we produce and compare our method against the conventional LERP method, as well as three state-of-the-art interpolation models trained on the original CMU MoCap dataset,. Specifically, we measure the performance of the RNN-based $\text{TG}_{\text{complete}}$ model [18], the BERT-based motion interpolation adaptation [12], and a transformer-based encoder-decoder approach [33]. We additionally provide results on a variant of our method trained without our RPA strategy. The experimental configuration for each state-of-the-art model is identical to their original implementations. To measure interpolation accuracy, we employ the standard $\ell_{2}$ positional distance (L2P), $\ell_{2}$ quaternion difference (L2Q), and NPSS [16] for visual similarity evaluation [18].
@@ 4.4 Cross-Skeleton Motion Re-Targeting Experiments
Table 2: Average unpaired LaFAN1 to CMU motion re-targeting performance of the primal skeleton method and our point cloud method. $\mathcal{L}_{\text{KNN}}$ is measured using 1024-point clouds at $K=8$. Independently sampled point clouds from the original LaFAN1 skeleton are provided as optimal $\mathcal{L}_{\text{KNN}}$ reference. All values are scaled by ×$\times 100$.
@@
Since our approach is largely focused on enabling non-native motion interpolation supervision, we additionally compare our PC-MRL method against a CITL model supervised by PS. Due to the suboptimal re-targeting performance of PS, models supervised by PS exhibit expectedly poor accuracy, as shown in Table 3.

### v2

@@ 2.1 Data-Driven Motion Modelling Methods
Contemporary research in motion modelling predominantly explore learned data-driven methods, using neural networks to capture, model, and establish correlations between the motion features of various skeletal configurations [35, 51, 55, 41, 29]. Motion features alone have demonstrated considerable efficacy in cross-skeleton imitation using a variety of strategies, including task-driven objectives [45, 19], diffusion-based generation [49, 25, 52], and latent feature consistency [2, 1, 46, 3, 7]. Latent consistency in particular learns inter-skeleton correlations without the need for paired datasets, which reduces its barrier of entry.
@@ 3.1 Quaternion-based Motion Representations
[…]
@@ 3.3 First-Frame Offset Quaternions
Crucially, FOQ is a roll-invariant rotation representation, that is, FOQ remains identical regardless of a bone’s absolute roll position in the original data. Trivially, to adjust the rotational roll of a bone $\mathbf{v}_{n}$, we perform the quaternion multiplication for each frame: $\mathbf{q}_{n,t}\times\mathbf{r}$, where $\mathbf{r}$ is a roll quaternion. As a quaternion on the roll axis of $\mathbf{v}_{n}$, $\mathbf{r}$ is constrained within $\mathbf{r}=\alpha+\beta(x\mathbf{i}+y\mathbf{j}+z\mathbf{k})$, where $x$,$y$,$z$ are the 3D XYZ values of $\mathbf{T}_{n}$, and $\alpha,\beta$ are scalars that control the roll magnitude. Note that $\mathbf{r}$ is constant throughout the sequence. Therefore, we can deduce the roll-invariant nature of any constant $\mathbf{r}$:
@@ 3.4 Motion Interpolation Transformer and Rest Pose Augmentation
To leverage our point cloud-based representation learning for cross-skeleton interpolation, a transformer model - CITL [34] is adapted, allowing for an efficient interpolation pipeline with existing datasets to alternative skeletons. We introduce two key modifications to CITL and its training strategy, focusing on enhancing motion quality. First, quaternion predictions are substituted with FOQ predictions. This is inspired by the RNN-based method $\text{TG}_{complete}$ [18], which predicts quaternion offsets from the preceding frame rather than raw quaternions, thereby achieving greater generalisability and superior quality in interpolation.
@@
[…]
@@ 4.1 Datasets
  • Human3.6M [22, 6] comprises of 3.6 million motion frames. The motions captured in Human3.6M tend to be less dynamic, as they were recorded on a relatively compact 4x3m stage. Its native skeleton includes 5 spinal bones, 4 bones for each limb, and an additional finger bone for each hand.
  • CMU Mocap [10] is a diverse motion dataset encompassing 113 categories. It features a mix of static and dynamic motions, captured on a 3x8m stage. The skeleton is identical to that of Human3.6M, with the exception of the pelvis, which is divided into two hip bones and a lower back bone.
For PC-MRL, we designated the CMU MoCap skeleton as $S_{B}$. Its native dataset is exclusively utilized for testing purposes, while LaFAN1 and Human3.6M are used simultaneously for training. [^1] The authors Clinton Mo, Kun Hu and Zhiyong Wang are signatories of the dataset licenses and produced all experimental results of the paper. Meta Inc. was not granted access to the datasets.
@@ 4.2 Implementation Details
In our experiments, point cloud obfuscation adopted a 256-point setting for the sampling process with $\sigma=0.05$ and each point was characterized by a 64-element vector. To maintain consistency, we ensure that all bone offsets and motion positions are defined in metre units. For skeletal reconstruction, 4 Point Transformer [54] layers was utilized to construct the neural network. Within each layer, the feature size was doubled, and the size of the point set was downsampled by a factor of 4 through a farthest-first traversal approach, eventually resulting in a singular feature vector. In the final step, two linear layers with ReLU activations were utilized to transform the resulted feature vector in a frame-wise manner, producing the output skeletal representation. During training, we set $K=8$ for the KNN loss to optimize the network. Additionally, the use of unit quaternion values in practice relies on modulus division for constraint, a process that can potentially lead to exploding gradients. To address this, we introduced an objective function $\mathcal{L}_{q}$ that measures the $\ell_{2}$ norms’ difference between the raw quaternion output and the expected unit quaternion. All experiments were trained and evaluated on a single NVIDIA GeForce RTX 3090 GPU.
@@ 4.3 Motion Interpolation Comparison against Supervised Methods
To demonstrate the efficacy of our method, we produce and compare our method against the conventional LERP method, as well as three state-of-the-art interpolation models trained on the original CMU MoCap dataset. Specifically, we measure the performance of the RNN-based $\text{TG}_{\text{complete}}$ model [18], the BERT-based motion interpolation adaptation [12], and a transformer-based encoder-decoder approach [34]. We additionally provide results on a variant of our method trained without our RPA strategy. The experimental configuration for each state-of-the-art model is identical to their original implementations. To measure interpolation accuracy, we employ the standard $\ell_{2}$ positional distance (L2P), $\ell_{2}$ quaternion difference (L2Q), and NPSS [16] for visual similarity evaluation [18].
@@ 4.4 Cross-Skeleton Motion Re-Targeting Experiments
Table 2: Average unpaired LaFAN1 to CMU motion re-targeting performance of the primal skeleton method and our point cloud method. $\mathcal{L}_{\text{KNN}}$ is measured using 1024-point clouds at $K=8$. Independently sampled point clouds from the original LaFAN1 skeleton are provided as optimal $\mathcal{L}_{\text{KNN}}$ reference. All values are scaled by ×100$\times$ 100× 100.
@@
Since our approach is largely focused on enabling non-native motion interpolation supervision, we additionally compare our PC-MRL method against a CITL model supervised by PS in Table 3. Due to the suboptimal re-targeting performance of PS, such interpolation models exhibit expectedly poor accuracy.